Q4) 2 points A researcher in the University of Toronto claims that the average life span of mice can be extended by as much aPlease write the full derivation of the answers. Thank you!
Answer #1

H_{0}: \mu1 - \mu2 \geq 8

H_{1}: \mu1 - \mu2 < 8

The pooled variance (sp2) = ((n1 - 1)s1^2 + (n2 - 1)s2^2)/(n1 + n2 - 2) = (14 * (2.8)^2 + 9 * (3.2)^2)/(15 + 10 - 2) = 8.7791

The test statistic is

t = \frac{(\bar x1 - \bar x2) - (\mu1 - \mu2)}{\sqrt{\frac{s_{p}^2}{n1} + \frac{s_{p}^2}{n2}}}

  = \frac{(37.6 - 32.1) - 8}{\sqrt{\frac{8.7791}{15} + \frac{8.7791}{10}}}

= -2.07

df = 15 + 10 - 2 = 23

At alpha = 0.05, the critical value is -t0.05, 23 = -1.714

Since the test statistic value is less than the critical value (-2.07 < -1.714), so we should reject the null hypothesis.

At 0.05 significance level, there is sufficient evidence to conclude that the average life span of mice can be extended by as much as 8 months when the calorie in their diet are reduced by approximately 40% from the time they are weaned.